what is meant by PHASE, PHASE CHANGE & INTERFERENCE in case of light ray ?

Grade:12

1 Answers

AskiitianExpert Shine
14 years ago

Hi

Wave can be added together either constructively or destructively. The result of adding two waves of the same frequency depends on the value of the phase of the wave at the point in which the waves are added.
Electromagnetic waves are subject to interference. For two sources of electromagnetic waves to interfere

  • The sources must have the same frequency and polarization.
  • The sources must be coherent.
  • The superposition principle must apply.

In physics, interference is the addition (superposition) of two or more waves that result in a new wave pattern. Interference usually refers to the interaction of waves which are correlated or coherent with each other, either because they come from the same source or because they have the same or nearly the same frequency.

Two non-monochromatic waves are only fully coherent with each other if they both have exactly the same range of wavelengths and the same phase differences at each of the constituent wavelengths.

The total phase difference is derived from the sum of both the path difference and the initial phase difference (if the waves are generated from two or more different sources). It can then be concluded whether the waves reaching a point are in phase (constructive interference) or out of phase (destructive interference).

 

The phase of an oscillation or wave is the fraction of a complete cycle corresponding to an offset in the displacement from a specified reference point at time t = 0. Phase is a frequency domain or Fourier transform domain concept, and as such, can be readily understood in terms of simple harmonic motion. The same concept applies to wave motion, viewed either at a point in space over an interval of time or across an interval of space at a moment in time. Simple harmonic motion is a displacement that varies cyclically, as depicted to the right.

It is described by the formula:

x(t) = A\cdot \sin( 2 \pi f t + \theta ),\,

where A is the amplitude of oscillation, f is the frequency, t is the elapsed time, and θ is the phase of the oscillation. The phase determines or is determined by the initial displacement at time t = 0. A motion with frequency f has period T=\frac{1}{f}.